I have a tough word problem that I just can not seem to get correct. Can someone please help?
Find the length of side a in the right triangle, if b=8 and c=10.




Answer by algebrahouse.com(1659)

"I have a tough word problem that I just can not seem to get correct. Can someone please help?
Find the length of side a in the right triangle, if b=8 and c=10."

Pythagorean theorem is a² + b² = c²
where a and b are the legs and c is the hypotenuse

a² + 8² = 10² {substituted given sides into Pythagorean Theorem}
a² + 64 = 100 {evaluated}
a² = 36 {subtracted 64 from both sides}
a = 6 {took square root of both sides}
www.algebrahouse.com